# Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output

### v2

@@ 3.2. Data-Plane: Prompt-Based Attacks
We investigate state-of-the-art (SOTA) LLM defenses against prompt-based attacks on the data plane. To this end, we apply AdvBench (advbench2023emnlp), which provides a set of prompt-based attacks, to gpt-4o. Results show that the model’s defenses are highly effective, with an attack success rate of only 1.1%. As illustrated in Figure 3a, most of gpt-4o’s refusal responses are short and deterministic (e.g., variations of “I’m sorry”), accounting for roughly 94% of all outputs. This behavior can be attributed to external guardrails that terminate generation and replace it with a brief refusal message—a mechanism also observed and exploited by MasterKey (masterkey2024ndss) using a time-based SQL injection method.
Beyond this, 25 out of 520 attempts yielded longer but still benign refusal explanations, largely driven by the model’s internal alignment. In rare cases (6 out of 520), both the internal alignment and external guardrails failed, allowing gpt-4o to produce harmful content. Overall, these results suggest that modern LLM providers employ strong guardrails that make prompt-based data-plane attacks difficult to execute successfully.
@@ 3.3. Control-Plane: A New Attack Surface
In light of this observation, we first develop a intuitive Proof-of-Concept attack called EnumAttack (detailed in §§ 4.1). When launched against models like gpt-4o, EnumAttack successfully compromises all 520 test cases in AdvBench (advbench2023emnlp), achieving a 100% success rate (Figure 3c). While highly effective, EnumAttack is relatively “exposed” as the malicious strings are explicitly present in the grammar, making it vulnerable to straightforward grammar-level guardrails.
However, we find that the attack surface can be made significantly more stealthy. By further decoupling the malicious payload across space (splitting intent between the data plane prompt and the control plane grammar) and time (interleaving payloads across multiple asynchronous requests), an adversary can construct a more evasive attack that is difficult to detect or defend against, which we instantiate as DictAttack (detailed in §§ 4.2). This reveals that the constrained decoding mechanism itself inherently lacks the necessary cross-plane security coordination. Our findings demonstrate an urgent need to protect this new, largely ignored attack surface to secure the entire LLM generation pipeline.
@@ 4. Constrained Decoding Attacks
Building on the aforementioned observations, we introduce the Constrained Decoding Attack (CDA) as a novel vulnerability and new dimension in LLM security.
Unlike traditional jailbreaks that rely on specially crafted prompts, CDA exploits the control plane by manipulating grammar rules used to constrain model outputs during decoding. This creates a new attack surface: instead of attacking the prompt (data plane), an adversary embeds malicious intent directly into the grammar that governs structured output. In this way, CDA leverages the constrained decoding mechanism itself to bypass guardrails and induce harmful behavior.
@@ 4.1. EnumAttack: Direct Manipulation via JSON Schema
Figure 4. Illustration of EnumAttack, where the malicious intent is embedded in the enum property of the JSON Schema (control plane), while the prompt (data plane) remains benign.
In our investigation of the control-plane attack surface, we first identified that the enum property in JSON Schema—designed to restrict fields to a fixed set of constants—provides a direct mechanism for forcing arbitrary malicious strings into the LLM’s generation context. We instantiate this as EnumAttack. By defining a schema field with a single-item enum containing the malicious query, the attacker leverages the determinism of the grammar engine to override the model’s internal safety alignment. Concurrent to our study, independent informal observations (EnumAttackBlog) have also noted the potential for exploiting enum fields in proprietary models, but lacked a systematic implementation or characterization of the underlying vulnerability.
@@
Chain EnumAttack. The EnumAttack illustrates how CDAs exploit shallow safety alignment (deepalignment2025iclr) to induce immediate harmful outputs. Building on this, we introduce Chain EnumAttack as a Proof-of-Concept to further reveal the fundamental vulnerability of LLM internal safety alignment when faced with user-forced content. By using a multi-stage strategy—where a model is first compromised via EnumAttack to generate harmful prefixes, which are then enforced as fixed grammar constraints in a subsequent turn—we demonstrate that even deeply aligned models can be coerced into completing harmful tasks when conditioned on such pre-filled malicious contexts. A detailed qualitative and quantitative analysis of how this process progressively breaks internal alignment is provided in §§ 5.4.
4.2. DictAttack: Benign Grammar + Prompt = Jailbreak
@@
During inference, the model receives the benign keys in the prompt and the dictionary in the grammar. The constrained decoding process then forces the model to select tokens that correspond to the values mapped to those keys in the dictionary. Neither the keys in the prompt nor the words in the dictionary are inherently harmful when viewed in isolation. Only when the model reconstructs the query by joining the dictionary values does the malicious intent emerge. This “Dual-Plane Decoupling” makes DictAttack exceptionally evasive, as a guardrail would need to semantically audit both planes simultaneously, potentially recalling all history context, to detect the threat. By increasing the synthesis ratio $k$, the attacker can dilute the ratio of harmful words in the dictionary, making detection via frequency analysis or semantic mismatch even more difficult.
Interleaved DictAttack. In practice, the attacker can further decouple DictAttack across time in multi-turn interactions. Specifically, the benign key sequence can be sent in an earlier turn (thus becoming part of the model’s conversational context / KV cache), while the dictionary-embedded schema is only provided in a subsequent request. This interleaving degrades many “dual-plane” auditing defenses that only inspect the current prompt+grammar pair, effectively reducing them back to single-plane checks. We refer to this variant as Interleaved DictAttack and evaluate its impact in §§§ 5.3.4.
@@ 4.3. Other CDAs
Beyond the attacks presented so far, other forms of CDAs are possible. For example, EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv), though designed as white-box methods that directly manipulate output logits, can also be realized indirectly via guided grammar.
APT (APT2025arxiv) leverages GuidedRegex to iteratively block refusal tokens, but the approach is non-functional in practice since its claimed backends, xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv), lack support for negative lookahead regex (e.g., “(?!Sorry)”). Even if supported, the method is inefficient, requiring exhaustive token-by-token trials.
Prompt-based jailbreaks remain orthogonal to our work and can be combined with CDAs. Template-based methods (MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), PAIR (PAIR2024arxiv), TAP (TAP2024nips)) can strengthen EnumAttack by crafting more complex “yes-prefix” fields, while linguistic (drattack2024emnlpfinding; puzzler2024aclfinding) and encoding attacks (zulu2024arxiv; base64attack2023nips; wei2023jailbroken) can further obscure malicious content by exploiting the mismatch between models’ advanced capabilities and their safety alignment.
@@ 5. Evaluation
AdvBench(advbench2023emnlp) 520 $\circ$ $\circ$
StrongREJECT(strongreject2024iclrws) 311 6 $\circ$
JailbreakBench(jailbreakbench2024nips) 100 10 $\circ$
HarmBench(harmbench2024arxiv) 100 3 $\circ$
@@
CDAs represent a shift in the jailbreaking paradigm by targeting the decoding process rather than relying on prompt engineering. By manipulating the structured output specifications through carefully crafted schema definitions (as demonstrated in our EnumAttack implementation), these attacks force models to generate content that would otherwise be filtered by both external guardrails and internal alignment mechanisms. We have revealed the vulnerability within the grammar control plane through the proposed attacks (EnumAttack and DictAttack), where plain jailbreak questions can be embedded in the grammar to cause jailbreaks, while no guardrails currently in this vulnerable path.
In this section, we first build a LLM guardrail system to mitigate CDAs, which is shown in Figure 6. Following MasterKey (masterkey2024ndss), which already establish a LLM system with prompt-based guardrail, we further enrich a grammar guardrail to protect the control plane as mitigation strategy. We provide a systematic evaluation of EnumAttack and DictAttack across different models, benchmarks and guardrail choices.
@@ 5.1. Experimental Setup
Attack methods. Besides our proposed EnumAttack (§§ 4.1) and DictAttack (§§ 4.2), we evaluate five baseline data-plane attacks: IJP (jailbreak2024ccs), DrAttack (drattack2024emnlpfinding), Puzzler (puzzler2024aclfinding), Zulu (zulu2024arxiv), and Base64 (base64attack2023nips). For comparative analysis against SOTA data-plane jailbreaks, we also include PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr). For each benchmark, we follow the methodology described in §§ 4.1 and §§ 4.2, applying EnumAttack and DictAttack to the harmful query intents provided by the respective dataset.
Datasets. Following previous works  (align2023nips; GCG2023arxiv; enforcedecode2024acl; JBShield2025usenixsecurity), we evaluate the performance of constrained decoding based attacks (CDAs) in six well-known benchmarks, whose statistics are shown in Table 1.
@@
  • Industrial Guardrails: We deploy llama-guard-3-8b (llamaguard2023arxiv) and OpenAI Moderation API (openaimoderation2023aaai), which are widely used for filtering prompts and responses.
@@
  • SOTA Academic Defenses: We evaluate against recent JBShield (JBShield2025usenixsecurity) and SelfDefend (selfdefend2025security), two state-of-the-art data-plane defenses proved to be effective on known jailbreak attacks.
@@
  • Dual-Plane Guard: We adapt SelfDefend (selfdefend2025security) to jointly audit the prompt and grammar, instantiated with gpt-4o as the auditing LLM (i.e., 4o guarding 4o when evaluating gpt-4o).
@@
Large language models. We evaluate a broad spectrum of 13 LLMs with native structured output support, categorized by their scale and deployment scenarios. For proprietary models, we test the widely-deployed gpt-4o, gpt-4o-mini, and gemini-2.0-flash, alongside the latest flagship models gpt-5 and gemini-2.5-pro. For open-weight models, we evaluate a range of architectures including llama-3.1-8b (llama3-2024website), qwen-2.5-32b (qwen2.5_2024arxiv), mistral nemo (mistralnemo2024web), phi-3.5-moe (phi3safety2024arxiv), and gemma-2-9b (gemma2_2024arxiv), as well as high-capability models like deepseek-v3/r1 (deepseekv3-2025arxiv; deepseekr1-2025arxiv) and gpt-oss-120b (gpt-oss). Throughout Evaluation, we report results by attack scenario: EnumAttack is primarily evaluated on mainstream deployed systems (gpt-4o, gpt-4o-mini, gemini-2.0-flash), while DictAttack is highlighted on flagship and high-capability models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). All models [^1] Checkpoints used: gpt-4o-2024-0806, gpt-4o-mini-2024-0718, gemini-2.0-flash-001, gpt-5, gemini-2.5-pro, deepseek-v3-0324, deepseek-r1-0528, and gpt-oss-120b. For all open-weight models, we specifically evaluate their instruction-tuned versions (e.g., llama-3.1-8b-instruct, qwen-2.5-32b-instruct, gemma-2-9b-it) to ensure standard safety alignment is active. We also exclude preview version LLMs due to their instability, like gemini-3.0-pro-preview. are evaluated using black-box API access to simulate realistic attack scenarios.
@@
Evaluation metrics. We first leverage the mainstream metric: Attack Success Rate (ASR), which quantifies the percentage of successful jailbreaks. ASR is a commonly used measure in LLM security research (GCG2023arxiv; align2023nips; enforcedecode2024acl; strongreject2024iclrws), typically relying on an external LLM to judge. Following this convention, we employ another powerful LLM gpt-4o (gpt4o2024web) as the judge in our evaluation, which is in line with previous work (strongreject2024iclrws). An attack is deemed successful only if the attacker’s query doesn’t trigger any safety flag during prompt and grammar auditing (i.e., bypassing guardrails), and the LLM itself answers the query as expected (i.e., bypassing internal alignment). Furthermore, for the guardrails deployed to defend against our attacks, we report their Detection Rate. This metric reflects the guardrail’s ability to identify malicious inputs; therefore, a lower detection rate indicates a more effective attack that successfully evades the defense mechanism.
@@
For local open-weight LLMs, we test both vllm 0.10.1 (vllm2023sosp) and sglang 0.4.10 (sglang2024arxiv) for serving. Both vllm and sglang support an OpenAI compatible server with structured output features supported by a grammar backend like xgrammar (xgrammar2024arxiv) or outlines (outlines2023arxiv). We use the OpenAI compatible server to call the local LLMs so that the evaluation for local models is consistent with the evaluation for proprietary models. For large-scale open-weight models such as deepseek-v3/r1 and gpt-oss-120b, due to resource limits, we utilize APIs provided by OpenRouter (openrouter) to evaluate them in realistic production serving scenarios.
5.2. CDAs Can Effectively Bypass Prompt Guardrails
In this subsection, we focus on evaluating the effectiveness of CDAs against prompt guardrails, thus only the prompt guard of Figure 6 is enabled. We evaluate along three dimensions:

### v3

@@ 3.2. Data-plane: Prompt-Based Attacks
We first probe state-of-the-art LLM defenses against prompt-based attacks on the data plane by running AdvBench (Chen et al., 2022) against gpt-4o, which yields only 1.1% ASR. As Figure 3a shows, $\sim$94% of responses are short deterministic refusals (e.g., variants of “I’m sorry”) that the external guardrail substitutes for the model’s actual generation, a behaviour also exploited by MasterKey (Deng et al., 2024); another 25/520 responses are longer benign refusals driven by internal alignment, and only 6/520 leak harmful content. Modern proprietary APIs therefore make direct data-plane attacks largely ineffective — a baseline we revisit quantitatively across five benchmarks in Table 4.
@@ 3.3. Control-plane: A New Attack Surface
This observation motivates two proof-of-concept CDAs developed in § 4: the intuitive EnumAttack and the more evasive DictAttack, which decouples the malicious payload across both space (data plane vs. control plane) and time (asynchronous requests). Together they expose that constrained decoding lacks the cross-plane security coordination needed to defend the LLM generation pipeline.
@@ 4. Constrained Decoding Attacks
Building on the aforementioned observations, we introduce the Constrained Decoding Attack (CDA) as a novel vulnerability and new dimension in LLM security. Unlike traditional jailbreaks that rely on specially crafted prompts, CDA exploits the control plane by manipulating grammar rules used to constrain model outputs during decoding. This creates a new attack surface: instead of attacking the prompt (data plane), an adversary embeds malicious intent directly into the grammar that governs structured output. In this way, CDA leverages the constrained decoding mechanism itself to bypass guardrails and induce harmful behavior. Concretely, every CDA decomposes into two stages: (1) control-plane injection, where schema-enforced logit masking forces the output into an attacker-chosen state — not necessarily a short or visibly harmful prefix, but possibly a richer grammar-specified structure (e.g., a dictionary or templated skeleton) whose pieces are benign in isolation and only encode malicious intent once the model assembles them; and (2) model-driven semantic continuation, where the model’s own coherence bias completes the harmful intent on that trajectory.
@@ 4.1. EnumAttack: Direct Manipulation via JSON Schema
Figure 4. Illustration of EnumAttack, where malicious intent is hidden in the JSON Schema enum (control plane) while the prompt (data plane) stays benign.
In our investigation of the control-plane attack surface, we first identified that the enum property in JSON Schema—designed to restrict fields to a fixed set of constants—provides a direct mechanism for forcing arbitrary malicious strings into the LLM’s generation context. We instantiate this as EnumAttack. By defining a schema field with a single-item enum containing the malicious query, the attacker leverages the determinism of the grammar engine to override the model’s internal safety alignment. Concurrent to our study, independent informal observations (Priyanshu, 2024) have also noted the potential for exploiting enum fields in proprietary models, but lacked a systematic implementation or characterization of the underlying vulnerability.
@@
Chain EnumAttack. The EnumAttack illustrates how CDAs exploit shallow safety alignment (Qi et al., 2025) to induce immediate harmful outputs. Building on this, we introduce Chain EnumAttack as a Proof-of-Concept to further reveal the fundamental vulnerability of LLM internal safety alignment when faced with user-forced content. By using a multi-stage strategy—where a model is first compromised via EnumAttack to generate harmful prefixes, which are then enforced as fixed grammar constraints in a subsequent turn—we demonstrate that even deeply aligned models can be coerced into completing harmful tasks when conditioned on such pre-filled malicious contexts. A detailed qualitative and quantitative analysis of how this process progressively breaks internal alignment is provided in §§ 5.4.
@@ 4.2. DictAttack: Benign Grammar + Prompt = Jailbreak
During inference, the model receives the benign keys in the prompt and the dictionary in the grammar. The constrained decoding process then forces the model to select tokens that correspond to the values mapped to those keys in the dictionary. Neither the keys in the prompt nor the words in the dictionary are inherently harmful when viewed in isolation. Only when the model reconstructs the query by joining the dictionary values does the malicious intent emerge. This “Dual-Plane Decoupling” makes DictAttack exceptionally evasive, as a guardrail would need to semantically audit both planes simultaneously, potentially recalling all historical context, to detect the threat. By increasing the synthesis ratio $k$, the attacker can dilute the ratio of harmful words in the dictionary, making detection via frequency analysis or semantic mismatch even more difficult.
Interleaved DictAttack. In practice, DictAttack can be split across turns. A benign key sequence is sent first and stored in the model’s context, while the dictionary-based schema is supplied later. This breaks many dual-plane defenses that only inspect the current prompt–grammar pair, reducing them to single-plane checks. We call this Interleaved DictAttack and evaluate it in §§§ 5.3.5.
@@ 4.3. Other CDAs
Beyond the attacks presented so far, other forms of CDAs are possible. For example, EnDec (Zhang et al., 2024) and JailMine (Li et al., 2024b), though designed as white-box methods that directly manipulate output logits, can also be realized indirectly via guided grammar.
APT (Li et al., 2025b) leverages GuidedRegex to iteratively block refusal tokens, but the approach is non-functional in practice since its claimed backends, xgrammar (Dong et al., 2025) and outlines (Willard and Louf, 2023), lack support for negative lookahead regex (e.g., “(?!Sorry)”). Even if supported, the method is inefficient, requiring exhaustive token-by-token trials.
Prompt-based jailbreaks remain orthogonal to our work and can be combined with CDAs. Template-based methods (MasterKey (Deng et al., 2024), LLMFuzzer (Yu et al., 2024), PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024)) can strengthen EnumAttack by crafting more complex “yes-prefix” fields, while linguistic (Li et al., 2024a; Chang et al., 2024) and encoding attacks (Yong et al., 2024; Wei et al., 2023a, b) can further obscure malicious content by exploiting the mismatch between models’ advanced capabilities and their safety alignment.
@@ 5. Evaluation
AdvBench(Chen et al., 2022) 520 $\circ$ $\circ$
@@
SorryBench(Xie et al., 2025) 440 44 21
[…]
To evaluate CDAs in realistic deployments, we extend the prompt-guard system of MasterKey (Deng et al., 2024) with an additional grammar guard for the control plane (Figure 6), and systematically evaluate EnumAttack and DictAttack across models, benchmarks, and guardrail choices.
@@ 5.1. Experimental Setup
Attack methods. Besides our proposed EnumAttack (§§ 4.1) and DictAttack (§§ 4.2), we evaluate five baseline data-plane attacks: IJP (Shen et al., 2024), DrAttack (Li et al., 2024a), Puzzler (Chang et al., 2024), Zulu (Yong et al., 2024), and Base64 (Wei et al., 2023a). For comparative analysis against SOTA data-plane jailbreaks, we also include PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024), and AutoDAN-Turbo (Liu et al., 2025). For each benchmark, we follow the methodology described in §§ 4.1 and §§ 4.2, applying EnumAttack and DictAttack to the harmful query intents provided by the respective dataset.
Datasets. Following previous works  (Carlini et al., 2023; Zou et al., 2023; Zhang et al., 2024, 2025), we evaluate the performance of constrained decoding based attacks (CDAs) in six well-known benchmarks, whose statistics are shown in Table 1.
@@
  • Industrial Guardrails: We deploy llama-guard-3-8b (AI, 2023a) and OpenAI Moderation API (Markov et al., 2023), which are widely used for filtering prompts and responses.
@@
  • SOTA Academic Defenses: We evaluate against recent JBShield (Zhang et al., 2025) and SelfDefend (Wang et al., 2025), two state-of-the-art data-plane defenses proved to be effective on known jailbreak attacks.
@@
  • Dual-Plane Guard: We adapt SelfDefend (Wang et al., 2025) to jointly audit the prompt and grammar, instantiated with gpt-4o as the auditing LLM (i.e., 4o guarding 4o when evaluating gpt-4o).
@@
Large language models. We evaluate a broad spectrum of 13 LLMs with native structured output support, categorized by their scale and deployment scenarios. For proprietary models, we test the widely-deployed gpt-4o, gpt-4o-mini, and gemini-2.0-flash, alongside the latest flagship models gpt-5 and gemini-2.5-pro. For open-weight models, we evaluate a range of architectures including llama-3.1-8b (Meta AI, 2024), qwen-2.5-32b (Qwen, 2025), mistral nemo (Mistral AI team, 2024), phi-3.5-moe (Haider et al., 2024), and gemma-2-9b (Team, 2024), as well as high-capability models like deepseek-v3/r1 (DeepSeek-AI, 2025b, a) and gpt-oss-120b (OpenAI, 2025). We report results by attack scenario: EnumAttack is primarily evaluated on mainstream deployed systems (gpt-4o, gpt-4o-mini, gemini-2.0-flash), while DictAttack is highlighted on flagship and high-capability models, including reasoning models like gpt-5, gemini-2.5-pro and deepseek-r1. All models [^1] Checkpoints used: gpt-4o-2024-0806, gpt-4o-mini-2024-0718, gemini-2.0-flash-001, gpt-5, gemini-2.5-pro, deepseek-v3-0324, deepseek-r1-0528, and gpt-oss-120b. For all open-weight models, we specifically evaluate their instruction-tuned versions (e.g., llama-3.1-8b-instruct, qwen-2.5-32b-instruct, gemma-2-9b-it) to ensure standard safety alignment is active. We also exclude preview version LLMs due to their instability, like gemini-3.0-pro-preview. Anthropic models are not included because their public APIs do not currently expose a JSON-Schema-based constrained decoding interface for structured output, so the threat model of CDA does not directly apply via their official API. are evaluated using black-box API access to simulate realistic attack scenarios.
@@
Evaluation metrics. We first leverage the mainstream metric: Attack Success Rate (ASR), which quantifies the percentage of successful jailbreaks. ASR is a commonly used measure in LLM security research (Zou et al., 2023; Carlini et al., 2023; Zhang et al., 2024; Souly et al., 2024), typically relying on an external LLM to judge. Following this convention, we employ another powerful LLM gpt-4o (OpenAI, 2024a) as the judge in our evaluation, which is in line with previous work (Souly et al., 2024). An attack is deemed successful only if the attacker’s query doesn’t trigger any safety flag during prompt and grammar auditing (i.e., bypassing guardrails), and the LLM itself actually satisfies the malicious intent (producing operationally specific harmful content), not merely fails to refuse; labeled judging examples are provided in our supplementary artifact. Furthermore, for the guardrails deployed to defend against our attacks, we report their Detection Rate. This metric reflects the guardrail’s ability to identify malicious inputs; therefore, a lower detection rate indicates a more effective attack that successfully evades the defense mechanism. To assess jailbreak result quality beyond binary success, we additionally evaluate the StrongREJECT (Souly et al., 2024) composite score on a subset of attacks (§§§ 5.3.2), which combines ASR with how convincing and specific the resulting harmful answer is to an adversary.
@@
For local open-weight LLMs, we test both vllm 0.10.1 (Kwon et al., 2023) and sglang 0.4.10 (Zheng et al., 2024) for serving. Both vllm and sglang support an OpenAI compatible server with structured output features supported by a grammar backend like xgrammar (Dong et al., 2025) or outlines (Willard and Louf, 2023). We use the OpenAI compatible server to call the local LLMs so that the evaluation for local models is consistent with the evaluation for proprietary models. For large-scale open-weight models such as deepseek-v3/r1 and gpt-oss-120b, due to resource limits, we utilize APIs provided by OpenRouter (OpenRouter Team, 2023) to evaluate them in realistic production serving scenarios.
5.2. CDAs Open an Attack Surface Orthogonal to Data-Plane Jailbreaks
Existing prompt-based jailbreaks act on the data plane; CDAs act on the control plane. The two surfaces are independent, and most data-plane techniques — prompt decomposition (DrAttack, Puzzler), template optimisation (PAIR, TAP, AutoDAN-Turbo), encoding tricks (Zulu, Base64) — can be folded into a CDA’s data-plane payload to harden it further. We illustrate the surface gap by showing that SOTA data-plane defenses, which crush prompt-based jailbreaks, do nothing against CDAs. To establish this asymmetry, we keep only the prompt guard of Figure 6 active and evaluate along three dimensions: